### v1

@@ Abstract
In this article, we start by finding the equations of motion for the abelian D4 brane worldvolume. Assuming Hypothesis H that the 4-flux in M-Theory is flux quantized in a non-abelian cohomology theory called 4-cohomotopy, and the three-flux on the M5 brane worldvolume in (twisted) 3-cohomotopy, we compute the minimal model for the cyclification of the quaternionic Hopf fibration which encodes the Bianchi identities for the fluxes on the D4 brane worldvolume after double dimensional reduction. The two pictures can be mapped to each other, and thus at the rational level, we conjecture a non-abelian relative cohomology theory for the D4 brane, fibered over the 10d Type IIA spacetime fluxes.
@@
As a next step, we want to understand the M5 brane[16, 17, 18, 19, 20] in M theory as a system fibered over the spacetime M theory charges, or in better words, a relative non-abelian cohomology twisted by the 11d M theory spacetime charges. The sourced Bianchi identity on the M5 brane worldvolume reads $\mathrm{d}H_{3}=\phi^{\ast}G_{4}$ tells us that at least at the rational level, it is the minimal model of the 7-sphere relative to the 4-sphere. The integral lift of this M5 brane story can be taken to be in the quaternionic Hopf fibration $h_{\mathbb{H}}:S^{7}\rightarrow S^{4}$ with homotopy fiber being $S^{3}$. Thus the M5 brane is said to be quantized in relative 3-cohomotopy twisted by the background spacetime 4-cohomotopy charges $(c_{3},c_{6})$ denoted by $\pi^{3+\phi^{\ast}(c_{3},c_{6})}$, where $\phi:\Sigma\rightarrow X$ is the usual embedding map throughout our article where $\Sigma$ is the worldvolume and $X$ is the 11D M theory target spacetime[21, eq.22]. Twisted cohomotopy also implies (twisted) String structures on the M5 brane worldvolume[22], anomaly cancellation[23] and the WZ term[24].
Now, we want to see this for the D4 brane, fibered over the 10d spacetime Type IIA fluxes. We want to see this from the relative cohomotopy point of view which encodes the Bianchi identities for the M5 brane. Upon circle reduction, the M5 brane is expected to give rise to the D4 brane[25, 26, 27]. From the previous discussion, it is expected that the Bianchi identities on the D4 brane should be fibered over the spacetime Type IIA fluxes. We would expect the CE algebra of cyclification of (relative) 3-cohomotopy to produce this system of D4 brane upon double dimension reduction of the M5 brane in 11d M theory.
@@
The structure of the paper is as follows. In section 2, we figure out the equations of motion [^1] We will interchangeably use the names Bianchi identity and the equation of motion. from the D4 brane abelian worldvolume actions. In section 3, we compute the Chevalley-Eilenberg algebra of the cyclification of $S^{7}$ relative to $S^{4}$, and we map the Bianchi identities to the ones derived from the worldvolume action. These two precisely can be mapped to each other, which motivates us to propose a rational relative cohomology for the system of D4 brane, fibered over the Type IIA String theory Bianchi identities, which we have takn to be rationally in $\mathrm{LS^{4}\sslash S^{1}}$, not in topological K theory twisted by the NS-NS three form flux. In section 4, we have made some concluding remarks.
@@ 2.1 M5 brane on a Circle
The abelian DBI Lagrangian density on the single D4 brane worldvolume reads
@@
We now write out the CS action for the D4 brane worldvolume here
@@
where $F_{4},F_{2}$ denote the spacetime RR fluxes corresponding to the RR potentials $C_{3},C_{1}$ in Type IIA String theory.
@@
Thus, to wrap things up, on the D4 brane worldvolume we have the Bianchi identities
@@
For CE algebras of loop spaces and cyclification, see [15, §3.2].
@@
Under the dgca morphism $\mathrm{CE}(\mathfrak{l}\mathrm{LS^{4}\sslash S^{1}})\rightarrow\mathrm{\Omega^{\bullet}_{\mathrm{dR}}(X^{1,9})}$ with $\omega_{2}\mapsto F_{2},g_{4}\mapsto F_{4},g_{3}\mapsto H_{3},g_{6}\mapsto F_{6},g_{7}\mapsto H_{7}$ we recover the usual Bianchi identities for 10d spacetime Type IIA theory fluxes
$\mathrm{d}F_{2}=0,\hskip 7.11317pt\mathrm{d}F_{4}=F_{2}\wedge H_{3},\hskip 7.11317pt\mathrm{d}F_{6}=F_{4}\wedge H_{3},\hskip 7.11317pt\mathrm{d}H_{3}=0,\hskip 7.11317pt\mathrm{d}H_{7}=-\tfrac{1}{2}F_{4}^{2}+F_{2}\wedge F_{6}.$
@@
These equations encode the Bianchi identities for the Type IIA spacetime RR fluxes in even degrees till degree 6, twisted by the NS-NS three-form flux. The last equation encodes the Chern-Simons term for the NS5 brane [15, page 12].
@@
: This CE algebra is ‘almost’ the same as $\mathrm{CE}(\mathfrak{l}(\mathrm{KU_{0}\sslash BU(1)}))$ except the fact that the $g_{7}$ does not arise in the twisted K theory.
$\begin{array}[]{l}\mathrm{CE}(\mathfrak{l}(\mathrm{KU_{0}\sslash BU(1)}))=\Big(\Lambda^{\bullet}(g_{2},g_{4},g_{6},h_{3};\mathrm{d});\mathrm{d}g_{4}=g_{2}h_{3},\\ \mathrm{d}g_{6}=h_{3}g_{4},\mathrm{d}g_{2}=0,\mathrm{d}h_{3}=0\Big).\end{array}$ (26)
@@
$\mathrm{CE}(\mathfrak{l}(\mathrm{L}S^{4}\sslash S^{1}))/\langle g_{7}\rangle\cong\mathrm{CE}(\mathfrak{l}(\mathrm{KU_{0}\sslash BU(1)})).$
@@
Now for the single M5 brane, we have the relative CE algebra
@@
$\begin{array}[]{l}\mathrm{CE}(\mathfrak{l}(\mathrm{L}S^{7}_{/S^{4}}\sslash S^{1}))=\Big(\Lambda^{\bullet}(\omega_{2},h_{3},g_{4},g_{7},h_{2},g_{3},g_{6};\mathrm{d}_{\mathrm{cyc}});\mathrm{d}_{\mathrm{cyc}}g_{4}=\omega_{2}g_{3},\mathrm{d}_{\mathrm{cyc}}g_{7}=-\tfrac{1}{2}g_{4}^{2}+\omega_{2}g_{6},\\ \mathrm{d}_{\mathrm{cyc}}h_{3}=g_{4}+\omega_{2}h_{2},\mathrm{d}_{\mathrm{cyc}}g_{3}=0,\mathrm{d}_{\mathrm{cyc}}g_{6}=g_{3}g_{4},\mathrm{d}_{\mathrm{cyc}}h_{2}=-g_{3},\mathrm{d}_{\mathrm{cyc}}\omega_{2}=0,\Big).\end{array}$ (29)
@@
In this brief note, we have derived in section 2 the equation of motion for the abelian D4 brane worldvolume fluxes in the presence of the Type IIA spacetime fluxes. In section 3, we have assumed Hypothesis H that the M5 brane is quantized in the quaternionic Hopf fibration, and from there we have used the corresponding CE algebra of the cyclification of the quaternionic Hopf fibration to generate the Type IIA spacetime fluxes and the Bianchi identities for the D4 brane worldvolume fluxes at the same time. The analysis exactly seems to match the one we have derived from the abelian worldvolume action for the D4 brane. Thus, we have closed the section by proposing a rational relative non-abelian cohomology for the D4 brane, fibered over the Type IIA 10d spacetime fluxes. It is worth mentioning here that we have that the rational non-abelian cohomology for the spacetime Type IIA fluxes in $\mathrm{LS^{4}}\sslash S^{1}$ , and not in $\mathrm{KU_{0}\sslash BU(1)}$, and then we have fibered the D4 brane fluxes over this as a relative cohomology. It would be interesting to understand it’s integral lift.

### v2

@@ Abstract
In this article, we start by re-deriving the equations of motion and Bianchi identities for the abelian D4 brane worldvolume. Noting that the 4-flux in M-Theory is rationally flux quantized in a non-abelian cohomology theory called 4-cohomotopy, and the three-flux on the M5 brane worldvolume in rational (twisted) 3-cohomotopy, we compute the minimal model for the cyclification of the quaternionic Hopf fibration which encodes the Bianchi identities for the fluxes on the D4 brane worldvolume after double dimensional reduction. The two pictures can be mapped to each other, and thus at the rational level, we conjecture a non-abelian relative cohomology theory for the D4 brane, fibered over the 10d Type IIA spacetime fluxes.
@@
As a next step, we want to understand the M5 brane[16, 17, 18, 19, 20] in M theory as a system fibered over the spacetime M theory charges, or in better words, a relative non-abelian cohomology twisted by the 11d M theory spacetime charges. The sourced Bianchi identity on the M5 brane worldvolume reads $\mathrm{d}H_{3}=\phi^{\ast}G_{4}$ tells us that at least at the rational level, it is the minimal model of the 7-sphere relative to the 4-sphere. The integral lift of this M5 brane story can be taken to be in the quaternionic Hopf fibration $h_{\mathbb{H}}:S^{7}\rightarrow S^{4}$ with homotopy fiber being $S^{3}$. Thus the M5 brane is said to be quantized in relative 3-cohomotopy twisted by the background spacetime 4-cohomotopy charges $(c_{3},c_{6})$ denoted by $\pi^{3+\phi^{\ast}(c_{3},c_{6})}$, where $\phi:\Sigma\rightarrow X$ is the usual embedding map throughout our article where $\Sigma$ is the worldvolume and $X$ is the 11D M theory target spacetime[21, eq.22]. Twisted cohomotopy also implies (twisted) String structures on the M5 brane worldvolume[22], anomaly cancellation[23] and the WZ term[24]. Also, the super exceptional embedding of the M5 brane and relation to 5d Super Yang Mills have been discussed in [25]. For this paper, the rational picture of the quaternionic Hopf fibration will suffice.
Now, we want to see this for the D4 brane, fibered over the 10d spacetime Type IIA fluxes. We want to see this from the relative cohomotopy point of view which encodes the Bianchi identities for the M5 brane. Upon circle reduction, the M5 brane is expected to give rise to the D4 brane[26, 27, 28, 29, 30]. From the previous discussion, it is expected that the Bianchi identities on the D4 brane should be fibered over the spacetime Type IIA fluxes. We would expect the CE algebra of cyclification of (relative) 3-cohomotopy to produce this system of D4 brane upon double dimension reduction of the M5 brane in 11d M theory. For the relationship between cyclification and double-dimensional reduction, see [31, §3].
@@
The structure of the paper is as follows. In section 2, we rederive the equations of motion [^1] We will interchangeably use the names Bianchi identity and the equation of motion. from the D4 brane abelian worldvolume actions. In section 3, we compute the Chevalley-Eilenberg algebra of the cyclification of $S^{7}$ relative to $S^{4}$, and we map the Bianchi identities to the ones derived from the worldvolume action. These two precisely can be mapped to each other, which motivates us to propose a rational relative cohomology for the system of D4 brane, fibered over the Type IIA String theory Bianchi identities, which we have takn to be rationally in $\mathrm{LS^{4}\sslash S^{1}}$, not in topological K theory twisted by the NS-NS three form flux. In section 4, we have made some concluding remarks.
@@ 2.1 M5 brane on a Circle
The abelian DBI Lagrangian density [29, eqns. 8, 60] on the single D4 brane worldvolume reads
@@
We now write out the CS action [29, eq. 60] for the D4 brane worldvolume here
@@
where $F_{4},F_{2}$ denote the spacetime RR fluxes corresponding to the RR potentials $C_{3},C_{1}$ in Type IIA String theory, and $F_{4}=dC_{3}-C_{1}\wedge H_{3}$.
@@
Thus, to wrap things up, on the D4 brane worldvolume we have the Bianchi identities (see also [26, eqns. 19,20])
@@
For CE algebras of loop spaces and cyclification, see [15, §3.2],[31, §3].
@@
Under the dgca morphism $\mathrm{CE}(\mathfrak{l}\mathrm{LS^{4}\sslash S^{1}})\rightarrow\mathrm{\Omega^{\bullet}_{\mathrm{dR}}(X^{1,9})}$ with $\omega_{2}\mapsto F_{2},g_{4}\mapsto F_{4},g_{3}\mapsto-H_{3},g_{6}\mapsto F_{6},g_{7}\mapsto H_{7}$ we recover the usual Bianchi identities for 10d spacetime Type IIA theory fluxes
$\mathrm{d}F_{2}=0,\hskip 7.11317pt\mathrm{d}F_{4}=-F_{2}\wedge H_{3},\hskip 7.11317pt\mathrm{d}F_{6}=-F_{4}\wedge H_{3},\hskip 7.11317pt\mathrm{d}H_{3}=0,\hskip 7.11317pt\mathrm{d}H_{7}=-\tfrac{1}{2}F_{4}^{2}+F_{2}\wedge F_{6}.$
@@
These equations encode the Bianchi identities for the Type IIA spacetime RR fluxes in even degrees till degree 6, twisted by the NS-NS three-form flux. The last equation encodes the Chern-Simons term for the NS5 brane [15, page 12] [^2] Our equation for $F_{6}$ differs from [15] by a factor of $\frac{1}{2}$..
@@
: This CE algebra is ‘almost’ the same as $\mathrm{CE}(\mathfrak{l}(\mathrm{KU_{0}\sslash BU(1)}))_{\leq 6}$ except the fact that the $g_{7}$ does not arise in the (6-truncated) twisted K theory. For this CE algebra, we have (see [15, §4.10]):
$\begin{array}[]{l}\mathrm{CE}(\mathfrak{l}(\mathrm{KU_{0}\sslash BU(1)}))_{\leq 6}=\Big(\Lambda^{\bullet}(g_{2},g_{4},g_{6},h_{3};\mathrm{d});\mathrm{d}g_{4}=g_{2}h_{3},\\ \mathrm{d}g_{6}=h_{3}g_{4},\mathrm{d}g_{2}=0,\mathrm{d}h_{3}=0\Big).\end{array}$ (26)
@@
$\mathrm{CE}(\mathfrak{l}(\mathrm{L}S^{4}\sslash S^{1}))/\langle g_{7}\rangle\cong\mathrm{CE}(\mathfrak{l}(\mathrm{KU_{0}\sslash BU(1)}))_{\leq 6}.$
@@
Now for the single M5 brane, we have the relative CE algebra (see [21, eq. 18])
@@
This notation $\mathrm{CE}(\mathfrak{l}\mathrm{L}S^{7}_{/S^{4}})$ means the minimal model of (the rational homotopy type of) $\mathrm{LS^{7}}$ relative to $\mathrm{LS^{4}}$ corresponding to the quaternionic Hopf fibration.
@@
$\begin{array}[]{l}\mathrm{CE}(\mathfrak{l}(\mathrm{L}S^{7}_{/S^{4}}\sslash S^{1}))=\Big(\Lambda^{\bullet}(\omega_{2},h_{3},g_{4},g_{7},h_{2},g_{3},g_{6};\mathrm{d}_{\mathrm{cyc}});\mathrm{d}_{\mathrm{cyc}}g_{4}=\omega_{2}g_{3},\mathrm{d}_{\mathrm{cyc}}g_{7}=-\tfrac{1}{2}g_{4}^{2}+\omega_{2}g_{6},\\ \mathrm{d}_{\mathrm{cyc}}h_{3}=g_{4}+\omega_{2}h_{2},\mathrm{d}_{\mathrm{cyc}}g_{3}=0,\mathrm{d}_{\mathrm{cyc}}g_{6}=g_{3}g_{4},\mathrm{d}_{\mathrm{cyc}}h_{2}=-g_{3},\mathrm{d}_{\mathrm{cyc}}\omega_{2}=0\Big).\end{array}$ (29)
@@
In this brief note, we have re-derived in section 2 the equation of motion for the abelian D4 brane worldvolume fluxes in the presence of the Type IIA spacetime fluxes. In section 3, we have taken into fact that the M5 brane is quantized rationally in the quaternionic Hopf fibration, and from there we have used the corresponding CE algebra of the cyclification of the quaternionic Hopf fibration to generate the Type IIA spacetime fluxes and the Bianchi identities for the D4 brane worldvolume fluxes at the same time. The analysis exactly seems to match the one we have derived from the abelian worldvolume action for the D4 brane. Thus, we have closed the section by proposing a rational relative non-abelian cohomology for the D4 brane, fibered over the Type IIA 10d spacetime fluxes. It is worth mentioning here that we have that the rational non-abelian cohomology for the spacetime Type IIA fluxes in $\mathrm{LS^{4}}\sslash S^{1}$ , and not in $\mathrm{KU_{0}\sslash BU(1)}$ , and then we have fibered the D4 brane fluxes over this as a relative cohomology. It would be interesting to understand it’s integral lift.